7. A 25% solution of alcohol is to mixed with a 40% solution to get 50 litres of a final mixture that is 30% alcohol. How much of each of the original solutions should be used?
Answer by checkley77(12844)

.40X+.25(50-X)=.30*50
.40X+12.5-.25X=15
.15X=15-12.5
.15X=2.5
X=2.5/.15
X=16.667 LITERS OF TH4 40% MIXTURE IS USED.
50=16.666=33.333 LITERS OF THE 25% MIXTURE IS USED.
PROOF:
.40*16.667+.25*33.333=.30*50
6.667+8.333=15
15=15